Badger World Wide, Inc., manufactures military, work, sport, and casual footwear and leather accessories under a variety of brand names, such as Moth, Lovlies, Walkers, and Bill Ladde. The following transactions occurred during a recent month. a. Made cash sales of $66,800 (example) b. Purchased $4,150 of additional supplies on account. Borrowed $71,600 on long-term notes. d. Purchased $20,200 in additional equipment, paying in cash e. Incurred $40,050 in selling expenses, paying two-thirds in cash and owing the rest on account. f Paid $7,500 in rent for this month, and also paid $7,500 for next month. Required: . For each of the transactions, complete the following table, indicating the account, amount, and direction of the effect+for increase and -for decrease) of each transaction under the accrual basis. Include revenues and expenses as subcategories of stockholders equity, as shown for the first transaction, which is provided as an example 2. Determine the companys preliminary net income.

Complete this question by entering your answers in the tabs below. Required1Required 2 For each of the transactions, complete the following table, indicating the account, amount, and direction of the effect (+ for increase and - for de of each transaction under the accrual basis. Include revenues and expenses as subcategories of stockholders equity, as shown for the first transa which is provided as an example. (Enter any decreases to account balances with a minus sign.) Transaction Assets Liabilities Stockholders Equity Cash 66,800 4,150 71,600 (20,200) (20,200) 40,050 7,500 (15,000) Sales Revenue 66,800 Accounts Payable 4,150 Cash Cash Equipment Cash Prepaid Rent Cash C. otes Payable (long-term)71,600 Accounts Payable

In the blue area that is slightly cut-out, it reads: For each of the transactions, complete the following table, indicating the account, amount, and direction of the effect (+ for increase and − for decrease) of each transaction under the accrual basis. Include revenues and expenses as subcategories of stockholders’ equity, as shown for the first transaction, which is provided as an example. (Enter any decreases to account balances with a minus sign.)
Answer #1

1. Badger Worldwide Inc.

Table of transaction:-

Transaction Assets Liabilities Stockholders Equity
Cash + Supplies + Equipment + Prepaid Rent = Accounts Payable + Notes Payable (long-term) + Revenue - Expenses
a. 66,800 + + + = + + 66,800 -
b. + 4,150 + + = 4,150 + + -
c. 71,600 + + + = + 71,600 + -
d. (20,200) + + 20,200 + = + + -
e. (26,700) + + + = 13,350 + + - 40,050
f. (15,000) + + + 7,500 = + + - 7,500
Balance 76,500 + 4,150 + 20,200 + 7,500 = 17,500 + 71,600 + 66,800 - 47,550

Calculations for transaction e.:-

Paid in Cash=(40,050×2/3)=$26,700

And remaining balance $13,350 ($40,050-$26,700) on Accounts.

2. Badger Worldwide Inc.

Income Statement

Items Amount Amount
Sales Revenue 66,800
Less:- Expenses
Selling Expenses 40,050
Rent Expenses 7,500
Total Expenses (47,550)
Net Income 19,250

Company's Preliminary Income is $19,250